OpenStudy (anonymous):

I need help with solving for (x+4)*x=15.5

OpenStudy (anonymous):

Hello, distribute the x to the terms in parenthesis, subtract 15.5 from both sides and then use the quadratic formula?

OpenStudy (anonymous):

I'm not sure what the quadratic formula. Is can you explain?

OpenStudy (anonymous):

\[ax^2 + bx + c = 0\] \[x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

Firstly get all of your values on one side and zero on the other

OpenStudy (anonymous):

I'm in 7th grade. Are you in high school? ...because you lost me. What is a, b, or c

OpenStudy (whpalmer4):

The quadratic formula gives you the solution for an equation \(ax^2+bx+c=0, a\ne0\).

OpenStudy (anonymous):

In my case which number is a, b, c

OpenStudy (whpalmer4):

\[(x+4)*x=15.5\]\[x^2+4x=15.5\]You could complete the square by taking half the coefficient of the \(x\) term, squaring it, and adding to both sides:\[x^2+4x+(\frac{4}{2})^2=15.5+(\frac{4}{2})^2\]Now rewrite the left hand side as a square:\[(x+2)^2=15.5+4=19.5\]Take the square root of both sides\[(x+2)=\pm\sqrt{19.5}\approx \pm4.416\] Solve for x\[x=-2\pm4.416 \]\[x=2.416, x = -6.416\] \(a\) is the coefficient of the \(x^2\) term, 1 in your case. \(b\) is the coefficient of the \(x\) term, 4 in your case. \(c\) I hope you can figure out :-)

OpenStudy (whpalmer4):

You have to put your equation in the right form before you can apply the formula. In this case that means distributing \(x\) as in the 2nd line of my post where I solve the equation.

OpenStudy (whpalmer4):

And if using the quadratic formula, you would want the equation to read \[x^2+4x-15.5=0\]giving \(a=1, b=4, c=-15.5\).

OpenStudy (anonymous):

Thank You